OpenStudy (anonymous):

Find the volume of the solid enclosed by the parabolic cylinder y=x^2 and the planes z=3y, z=2+y

OpenStudy (anonymous):

Would it be: \[\int\limits_{}^{}\int\limits_{R}^{}(2+y)-3ydydx?\]But I'm not sure about the limits of integration and where y=x^2 comes in

OpenStudy (turingtest):



OpenStudy (anonymous):

Ok, so one bound is 1 to 3! What about the other?

OpenStudy (turingtest):



OpenStudy (anonymous):

But, if I just needed to do that, then I would be using x^2 at all?

OpenStudy (anonymous):

*wouldn't

OpenStudy (turingtest):

you still gotta integrate that over D, make it a function of y

OpenStudy (anonymous):

What do you mean? Haven't we found the bounds of integration and the function?

OpenStudy (turingtest):

what function are you using?

OpenStudy (anonymous):

2+y-3y? This problem is one from a section and the instructions at the top of the exercise list say to find the volume of the solid by subtracting the two volumes.

OpenStudy (turingtest):

the way I figured it the function we would use here would be sqrt(y)... or maybe 2sqrt(y) to make up for the - part I'm actually a bit unsure

OpenStudy (turingtest):

the first one seems to be your exact problem

OpenStudy (anonymous):

Could you explain how they got the limits for those?

OpenStudy (turingtest):

I don't understand the dA=2xdy part

OpenStudy (anonymous):

Ha, I'm not sure I understand much of it either!